OpenStudy (anonymous):

how do i find the surface area of a right cone? with the measurements of 5ft as the radus and 12ft as a hight

OpenStudy (amistre64):

do we include the base area?

OpenStudy (amistre64):

if you cut the cone up its length and splay it out; we get a part of a circle; with radius 12 and arc length of 5

OpenStudy (amistre64):

the total area of the circle would be 2pi(12) = 24pi 24pi/5 is the length of the arc area of a sector is: (radian angle)r^2/2

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

arc length = angle(r) 5 = angle(12) angle = 5/12 ; in radians 5(12)^2 area = ------- = 5(6) = 30 if I did it right :) 12*2

OpenStudy (amistre64):

if that includes that area of the base then add on another 25pi

OpenStudy (anonymous):

thanks